OpenStudy (anonymous):

How much would $200 invested at 7% interest compounded annually be worth after 5 years?

OpenStudy (anonymous):

A=P(1+r)^years of compounding A=(1+7%)^5

OpenStudy (anonymous):

A=200(1+0.07)^ 5